Okay I have 3 questions that if you can solve them and show the work please!!!
1) log(x+2)+log(x-1)=1
2) write and exponential function whose graph passes through (2,6) and (6,8).
30 Write a power function whose graph passes through (2,4)and (6,8).
Please as soon as possible.
Thank you!

Answer by user_dude2008(1862)

1)

log(x+2)+log(x-1)=1

log[(x+2)(x-1)]=1

(x+2)(x-1)=10^1

(x+2)(x-1)=10

x^2+x-2=10

x^2+x-12=0

(x+4)(x-3)=0

x+4=0 or x-3=0

x=-4 (ignore) or x=3


only answer is x=3



2)

y=a*b^x

6=a*b^2

a=6/b^2

y=a*b^x

8=(6/b^2)*b^6

8=6b^4

b=(8/6)^(1/4)

b = 1.07456993

a=6/b^2

a=6/(1.07456993)^2


a=5.19615244

equation is y = 5.19615244*1.07456993^x



3)

Is equation supposed to be linear, quadratic, cubic, etc?